Super Systems Ltd. manufactures a number of computer hardware products. The following variable cost standards have been set for one of their products.
Particulars Standard Price/rate
Direct materials $10 per pound
Direct labor ?
Variable manufacturing overhead $5.375 per hour
Total standard variable cost to manufacture the product is $75 per product.

During February, 850 direct labor hours were worked and 600 units of the product were manufactured. Each unit of product required one unit of raw material. The actual cost per unit manufactured during February was $0.30 higher than the standard cost.

For the month of February, the standard cost of production for direct labor was $14,000 and for variable manufacturing overhead was $4,300. Actual variable manufacturing overhead was $4,340. Actual costs incurred for materials used were $24,000. The material quantity usage variance was an unfavorable $700. Fixed overhead per direct labor is $3.5 (assuming normal capacity of 900 labor hours). Actual fixed overhead incurred was $3,500 (use labor hours as cost driver for overhead costs).

There was no stock of materials in the beginning of the February however, at the end of the month there were 100 units of materials.

Question 2 - Calculate the Standard quantity of materials per unit of product

$3.50

$6.88

$3.05

$4.45
Answer #1

Standard variable manufacturing overhead cost = $4,300

Standard variable manufacturing overhead per hour = $5.375

Standard labor hours for producing 600 units = 4340/5.375 = 800 hours

Standard direct labor cost = $14,000

Standard direct labor cost per hour = 14000/800 = $17.5 per hour

Standard labor hours to produce one unit = 800 hours/600 units = 1.33 hours

Calculation of standard quantity of material per unit of product:

Standard direct labor cost per unit of product = 1.33 hours * $17.5 = $23.275

Standard Variable manufacturing overhead cost = 1.33 hours * $5.375 = $7.14875

Total standard variable cost per product = $75

Standard direct material cost per unit of product = $75-($7.14875+$23.275) = $44.57625

Standard direct material cost per pound = $10

Standard quantity of materials per unit of product = 44.57625/10 = 4.45 pounds.

Thanks,

Let me know if you have any doubts.